OpenStudy (anonymous):

An archer releases an arrow from a shoulder height of 1.39 m. When the arrow hits the target 18 m away, it hits point A. When the target is removed, the arrow lands 45 m away. Find the maximum height of the arrow along its parabolic path. Ive seen this question on this website but nobody ever came up with a clear answer ive been trying to figure it out for hours can somebody help??

OpenStudy (anonymous):

Anyone?? about to throw the towel in on this one

OpenStudy (anonymous):

Let's not throw the towel JUST YET haha

OpenStudy (anonymous):

I would say that you know the change in the y-value is -1.39m over 45m right?

OpenStudy (anonymous):

So then we know that...\(\Delta x=45m\) and also we know that if we have an initial speed at an angle, \(v_i\), it can be broken up into x, and y components:  With projectile motion, the x-component remains constant so then we know that: \[v_{ix}=v_i\cos\theta=\frac{\Delta x}{\Delta t}\] So then, therefore, we know that for a y-displacement of \(\Delta y\), a time of \(\Delta t\), an acceleration of \(a\), and an initial speed of \(v_i\), we have: \[\eqalign{ \Delta y=\ v_{iy}\Delta t+\frac{a}{2}\Delta t^2 }\]

OpenStudy (anonymous):

would you like to see what i have so far? this is confusing and very frustrating

OpenStudy (anonymous):

Sure! Let's see! I admit, it is NOT a very easy question haha

OpenStudy (anonymous):

y=f(x)=ax^2+bx+c f(0)=1.39 f(18)=h f(45)=0 a(0)^2+b(0)+c=1.39 a(18)^2+b(18)+c=h a(45)^2+b(45)+c=0 324a+18b=h-1.39 2025a+45b=-1.39

OpenStudy (anonymous):

So you have three equations: \[\eqalign{ &a(0)^2+b(0)+c=1.39\to c=1.39\\ &a(18)^2+b(18)+c=h\to324a+18b+1.39=h \\ &a(45)^2+b(45)+c=0\to2024a+18b+1.39=0 \\ &\\ &\\ &\\ &324a+18b+1.39=h\to a=\frac{h-1.39-18b}{324}\\ &2024a+45b+1.39=0\to a=-\frac{(45b+1.39)}{2024}\\ &\\&\\&\\ &\frac{h-1.39-18b}{324}=-\frac{(45b+1.39)}{2024} \\ &2024(h-1.39-18b)=-324(45b+1.39) \\ &2024h-2813.36-36432b=-14580b-450.36 \\ &\\ &h=\frac{21852b+2363}{2024} \\ }\]

OpenStudy (anonymous):

So that is the max height of its trajectory?

OpenStudy (anonymous):

well, as a function of the \(b\) value, yes, but otherwise, not a complete number

OpenStudy (anonymous):

Err I got to go! I'll take a look at this in more depth later!

OpenStudy (anonymous):

someone help